# SplatFormer: Point Transformer for Robust 3D Gaussian Splatting

### v2

@@ Appendix E Comparisons with Baselines
Figure E.1: Failure Case. While our method effectively reduces artifacts in 3DGS (Kerbl et al., 2023) and outperforms SplatFields (Mihajlovic et al., 2024), it does not fully restore some high-frequency details. MipNeRF360 (Barron et al., 2022) excels in detail modeling but suffers from floating issues.
Appendix F Limitations and Future Directions
@@
Fine-grained Details. Our approach occasionally struggles to recover high-frequency details, particularly in complex textures, as shown in Fig. E.1. While our method reduces artifacts in 3DGS and outperforms previous variants like SplatFields, it still requires improvement in rendering texture details. The limitation may stem from the restricted capacity of our current point transformer backbone which uses grid pooling on the input point cloud to expand the receptive field. Larger grid resolutions and smaller pooling strides can be beneficial, but this requires more computational budgets. Future work can innovate the design of the point transformer architecture, e.g. integrating multi-resolution hierarchy, to capture and recover high-frequency details. Designing a trainable adaptive population mechanisim to densify Gaussians in high-frequency regions can also help represent the details.
@@
Table F.1: Results on MVImgNet.
@@
Unbounded Scenes and Diverse Camera Setups. While this paper primarily focuses on object-centric scenes in this project, the concept of learning data-driven priors for out-of-distribution views holds promise for extending to unbounded, in-the-wild scenes. Another promising direction is enhancing OOD renderings across more diverse camera setups beyond the bottom-to-top task. To be more specific, MVImgNet (Yu et al., 2023) serves as a strong testbed, offering substantial real-world multi-view images. We also conduct some preliminary experiments to demonstrate its potential. In MVImgNet, each scene is captured via a semi-circular camera trajectory around an object. To study OOD renderings, we split each trajectory into frontal and side views, using one as input and the other as test views. In these OOD test views, 3DGS produces significant artifacts. To adapt our method to this OOD-NVS setting, we constructed a training set with 4k MVImgNet scenes featuring flawed 3DGSs and multi-view images, and trained SplatFormer accordingly. We also create a evaluation set consisting of 70 held-out scenes using the same front-side camera setup. For each evaluation scene, we use the trained SplatFormer to refine the flawed 3DGS, trained only on the frontal/side input views, in a single feed-forward pass. Then we render the refined 3DGS from the side/frontal deviated test views. The evaluation metrics are reported in Tab. F.1, and visual results are shown in Fig. F.1. SplatFormer can remove floaters in the empty space and achieves better metrics than 3DGS. However, it struggles to refine the geometry and appearance of foreground objects. This limitation may stem from the normalization and downpooling operations in the point transformer, which disproportionately downscales foreground objects compared to the large-scale background point clouds, making it difficult for SplatFormer to capture foreground details. We hypothesize that designing a novel adaptive downpooling mechanism within the point transformer may address this issue. Additionally, a divide-and-conquer strategy—i.e., decomposing scenes into objects and background and processing each component separately—could also be beneficial. We plan to explore these directions in future work.
Camera Setup 3DGS Ours GT
Refer to caption Refer to caption Refer to caption Refer to caption
Camera Setup 3DGS Ours GT
Refer to caption Refer to caption Refer to caption Refer to caption
Figure F.1: The potential of SplatFormer in handling in-the-wild, unbounded scenes. When trained on MVImgNet (in-the-wild unbounded scenes), SplatFormer learns to partially remove floaters, though refining objects’ flawed geometry remains a challenge.
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
@@
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
@@
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
@@
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
@@
Figure F.2: Results on ShapeNet-OOD. We compare our method with baselines: SyncDreamer (Liu et al., 2023b), LaRa (Chen et al., 2024a), SSDNeRF (Chen et al., 2023), 3DGS (Kerbl et al., 2023), Nerfbusters (Warburg* et al., 2023), SplatFields (Mihajlovic et al., 2024), InstantSplat Fan et al. (2024), 2DGS (Huang et al., 2024a), FSGS (Zhu et al., 2024), InstantNGP (Müller et al., 2022), and MipNeRF360 (Barron et al., 2022).
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
@@
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
@@
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
@@
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
Four of Inputs MipNeRF360 Nerfbusters InstantNGP SyncDreamer LaRa SSDNeRF
GT Ours 2DGS SplatFields FSGS 3DGS InstantSplat
Figure F.3: Results on Objaverse-OOD. We compare our method with baselines: SyncDreamer (Liu et al., 2023b), LaRa (Chen et al., 2024a), SSDNeRF (Chen et al., 2023), 3DGS (Kerbl et al., 2023), Nerfbusters (Warburg* et al., 2023), SplatFields (Mihajlovic et al., 2024), InstantSplat Fan et al. (2024), 2DGS (Huang et al., 2024a), FSGS (Zhu et al., 2024), InstantNGP (Müller et al., 2022), and MipNeRF360 (Barron et al., 2022).
@@
Figure F.4: Results on GSO-OOD. We compare SplatFormer, trained on Objaverse scenes, with Nerfbusters (Warburg* et al., 2023), 2DGS (Huang et al., 2024a) and MipNeRF360 (Barron et al., 2022).
@@
Figure F.5: Results on Real-World iPhone OOD. We compare SplatFormer trained on Objaverse with Nerfbusters (Warburg* et al., 2023), MipNeRF360 (Barron et al., 2022), 2DGS (Huang et al., 2024a), and 3DGS (Kerbl et al., 2023).

### v3

@@ Appendix G Baseline Implementations
Figure G.1: Failure Case. While our method effectively reduces artifacts in 3DGS (Kerbl et al., 2023) and outperforms SplatFields (Mihajlovic et al., 2024), it does not fully restore some high-frequency details. MipNeRF360 (Barron et al., 2022) excels in detail modeling but suffers from floating issues.
Appendix H Limitations and Future Directions
@@
Fine-grained Details. Our approach occasionally struggles to recover high-frequency details, particularly in complex textures, as shown in Fig. G.1. While our method reduces artifacts in 3DGS and outperforms previous variants like SplatFields, it still requires improvement in rendering texture details. The limitation may stem from the restricted capacity of our current point transformer backbone which uses grid pooling on the input point cloud to expand the receptive field. Larger grid resolutions and smaller pooling strides can be beneficial, but this requires more computational budgets. Future work can innovate the design of the point transformer architecture, e.g. integrating multi-resolution hierarchy, to capture and recover high-frequency details. Designing a trainable adaptive population mechanisim to densify Gaussians in high-frequency regions can also help represent the details.
@@
Table H.1: Result of different input view distributions. For 3DGS trained on high-elevation views views, our trained SplatFormer achieves only limited improvement in low-elevation test views.
PSNR SSIM LPIPS
3DGS 18.01 0.697 0.297
SplatFormer (0-shot) 18.22 0.712 0.283
Diverse Camera Setups. Our current model focuses on a specific input view distribution where the input views provide full angular coverage along one axis (e.g., azimuth) while having limited angular coverage along another axis (e.g., elevation). The model trained on our current OOD datataset struggles to enhance quality when faced with significantly different input view distributions. For example, in Objaverse-OOD test set, if high-elevation views ($\phi\geq 50\degree$) are used as input views and low-elevation ($\phi\leq 10\degree$) views are used as test views, our current SplatFormer model provides limited improvement in quality for low-elevation test views. We report the results in Tab. H.1. We attribute this limitation to two factors. First, high-elevation input views capture only a small portion of the scene, leaving much of the lower regions unobserved - areas typically covered by low-elevation views. Since SplatFormer does not learn to generate (or ’hallucinate’) completely unseen parts of the scene, it cannot correct artifacts in these unobserved regions. Second, variations in the distribution of input views produce different types of 3DGS artifacts, some of which differ significantly from those encountered during training. We believe this limitation can be addressed by creating a more diverse OOD synthetic training dataset with a wider range of input view trajectories. Equipping point transformer with generative capability and likelihood estimation can be another promising direction.
Table H.2: Results on MVImgNet.
@@
Unbounded Scenes. While this paper primarily focuses on object-centric scenes in this project, the concept of learning data-driven priors for out-of-distribution views holds promise for extending to unbounded, in-the-wild scenes. Since MVImgNet (Yu et al., 2023) serves as a strong testbed, offering substantial real-world multi-view images, we also conduct some preliminary experiments to demonstrate its potential. In MVImgNet, each scene is captured via a semi-circular camera trajectory around an object. To study OOD renderings, we split each trajectory into frontal and side views, using one as input and the other as test views. In these OOD test views, 3DGS produces significant artifacts. To adapt our method to this OOD-NVS setting, we constructed a training set with 4k MVImgNet scenes featuring flawed 3DGSs and multi-view images, and trained SplatFormer accordingly. We also create a evaluation set consisting of 70 held-out scenes using the same front-side camera setup. For each evaluation scene, we use the trained SplatFormer to refine the flawed 3DGS, trained only on the frontal/side input views, in a single feed-forward pass. Then we render the refined 3DGS from the side/frontal deviated test views. The evaluation metrics are reported in Tab. H.2, and visual results are shown in Fig. H.1. SplatFormer can remove floaters in the empty space and achieves better metrics than 3DGS. However, it struggles to refine the geometry and appearance of foreground objects. This limitation may stem from the normalization and downpooling operations in the point transformer, which disproportionately downscales foreground objects compared to the large-scale background point clouds, making it difficult for SplatFormer to capture foreground details. We hypothesize that designing a novel adaptive downpooling mechanism within the point transformer may address this issue. Additionally, a divide-and-conquer strategy—i.e., decomposing scenes into objects and background and processing each component separately—could also be beneficial. We plan to explore these directions in future work.
Camera Setup 3DGS Ours GT  Camera Setup 3DGS Ours GT
Figure H.1: The potential of SplatFormer in handling in-the-wild, unbounded scenes. When trained on MVImgNet (in-the-wild unbounded scenes), SplatFormer learns to partially remove floaters, though refining objects’ flawed geometry remains a challenge.
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Figure H.2: Results on ShapeNet-OOD. We compare our method with baselines: SyncDreamer (Liu et al., 2023b), LaRa (Chen et al., 2024a), SSDNeRF (Chen et al., 2023), 3DGS (Kerbl et al., 2023), Nerfbusters (Warburg et al., 2023), SplatFields (Mihajlovic et al., 2024), InstantSplat Fan et al. (2024a), 2DGS (Huang et al., 2024a), FSGS (Zhu et al., 2024), InstantNGP (Müller et al., 2022), and MipNeRF360 (Barron et al., 2022).
3DGS Ours GT      3DGS Ours GT
++++++RGB
++++++Depth
++++++Normal
3DGS Ours GT      3DGS Ours GT
++++++RGB
++++++Depth
++++++Normal
3DGS Ours GT      3DGS Ours GT
++++++RGB
++++++Depth
++++++Normal
Figure H.3: Geometry comparison. While 3DGS is not designed for accurate geometry reconstruction, our method can still enhance the rendered depth and normal of 3DGS.
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Four of Inputs GT MipNeRF360 Nerfbusters InstantNGP LaRa SSDNeRF SyncDreamer
Ours 2DGS SplatFields FSGS 3DGS InstantSplat EscherNet
@@
Figure H.4: Results on Objaverse-OOD. We compare our method with baselines: SyncDreamer (Liu et al., 2023b), EscherNet (Kong et al., 2024), LaRa (Chen et al., 2024a), SSDNeRF (Chen et al., 2023), 3DGS (Kerbl et al., 2023), Nerfbusters (Warburg et al., 2023), SplatFields (Mihajlovic et al., 2024), InstantSplat Fan et al. (2024a), 2DGS (Huang et al., 2024a), FSGS (Zhu et al., 2024), InstantNGP (Müller et al., 2022), and MipNeRF360 (Barron et al., 2022).
@@
Figure H.5: Results on GSO-OOD. We compare SplatFormer, trained on Objaverse scenes, with Nerfbusters (Warburg et al., 2023), 2DGS (Huang et al., 2024a) and MipNeRF360 (Barron et al., 2022).
@@
Figure H.6: Results on Real-World iPhone OOD. We compare SplatFormer trained on Objaverse with Nerfbusters (Warburg et al., 2023), MipNeRF360 (Barron et al., 2022), 2DGS (Huang et al., 2024a), and 3DGS (Kerbl et al., 2023).